OpenStudy (anonymous):

Suppose p varies directly as d, and p = 2 when d = 7. What is the value of d when p = 10?

OpenStudy (anonymous):

a. d=7/5 b. d= 20/7 c. d= 15 or d. d= 35?

OpenStudy (anonymous):

d= 35

OpenStudy (anonymous):

mk thx:)